# A pipeline for enabling path-specific causal fairness in observational health data

### v1

@@ 5.4.2. Systemic Lupus Erythematosus (SLE)
When examining the impact of gender on SLE, the NIE and SE are insignificant, and the NDE is small (0.002[0.002,0.001]-0.002\ [-0.002,-0.001]). The NDE and NIE increase significantly in the baseline model (NDE: 0.038[0.047,0.025]-0.038\ [-0.047,\ -0.025]; NIE: 0.012[0.021,0.008]-0.012\ [-0.021,\ -0.008]), with the NDE increasing more than the NIE. The negative NDE indicates that women are more likely to be diagnosed than men, consistent with the low diagnosis rates of SLE in men (Tan et al., 2012; Simard et al., 2022). The NIE suggests that individuals with a ”male” distribution are more likely to be diagnosed with SLE than individuals with data drawn from a ”female” distribution. Due to direct bias against men when diagnosing SLE and the larger size of the NDE in both the data and the baseline model, we focus on reducing the NDE for SLE.
5.4.3. Type 2 Diabetes Mellitus (T2DM)
@@
In the T2DM cohort, the dataset reflects a significant NDE (0.014[0.002,0.016]0.014\ [0.002,0.016]) that indicates overdiagnosis of T2DM in Black individuals, consistent with epidemiological literature (Deng et al., 2025; Hackl et al., 2024). The NDE is exacerbated in the baseline model (0.033[0.027,0.134]0.033\ [0.027,0.134]), which also worsens the NIE (0.059[0.053,0.107]0.059\ [0.053,0.107]). The NIE suggests that an individual with clinical features drawn from a ”white” population distribution is more likely to be diagnosed with T2DM; this is consistent with known algorithmic biases (Cronjé et al., 2023). The SE remains insignificant in the data and baseline model. Thus, we focus on jointly reducing the NDE and NIE for T2DM risk prediction.
5.4.4. Schizophrenia (SCZ)
@@
The NDE, NIE, and SE are all similar in magnitude (NDE: 0.018[0.005, 0.031]0.018\ [0.005,\ 0.031]; NIE: 0.017[0.003,0.027]0.017\ [0.003,0.027]; SE: 0.016[0.017,0.012]-0.016\ [-0.017,\ -0.012]) in the data. The positive NDE is reflective of overdiagnosis of SCZ in Black patients compared to white patients, while the positive NIE is reflective of the fact that an individual with clinical features drawn from a ”white” population distribution is more likely to be diagnosed with SCZ. The negative SE suggests that the confounder variables in $Z$ (gender, healthcare utilization) may play a larger role in diagnosis for Black patients than white patients. Of these three effects, the NDE reflects known overdiagnosis of SCZ in Black individuals (Trierweiler et al., 2000), while the NIE and SE may reflect broader disparities in quality of care (Fiscella and Sanders, 2016) and healthcare utilization (Manuel, 2018), respectively. In the baseline model, we observe a significant exacerbation in NDE (Model: 0.120[0.104, 0.134]0.120\ [0.104,\ 0.134]) and SE (Model: 0.069[0.073,0.061]-0.069\ [-0.073,-0.061]). The NIE remains similar the data (0.015[0, 0.039]0.015\ [0,\ 0.039]). These causal effects suggest that the model is more reliant on race than a clinician might be (NDE), and that the disparate weight on confounding variables (healthcare utilization, gender) is more pronounced in the model than in clinical practice (SE). Based on the tenfold increase in NDE between the raw data and the baseline model and the known problems with overdiagnosis of schizophrenia in Black individuals, we prioritize reducing the NDE in the SCZ risk prediction case. Additionally, we treat reduction of SE as a secondary target, as the effect may be due to disparities in healthcare utilization.
5.5. Perform and evaluate fairness interventions
@@
We discuss the efficacy and impact of all interventions outlined in Section 4.5 for each clinical prediction task. We present the AUROC-effect tradeoff for each of the ”target” effects (Figure 4).
@@
Figure 4. Tradeoffs between model performance and target path-specific effects across all tasks. We plot the target path-specific effects for each intervention (NDE, NIE, and/or SE) against the model performance (AUROC) for each of the ”target” effects (Section 4.4). (A) and (E) show the NDE and NIE for AMI; (B) shows the NDE for SLE; (C) and (F) show the NDE and NIE for T2DM; (D) and (G) show the NDE and SE for SCZ.
@@
4 columns by 2 rows of scatter plots. Points are colored differently based on the fairness intervention (included in the legend) and have associated error bars. The x-axis is the path-specific causal effect and the y-axis shows AUROC.
5.5.1. Acute Myocardial Infarction (AMI)
@@
As shown in Figure 4A and 4E, the baseline model and demographic-unaware model are associated with the highest model performance (Baseline AUROC: 0.812[0.796, 0.825]0.812\ [0.796,\ 0.825]; demographic unawareness: 0.814[0.799, 0.832]0.814\ [0.799,\ 0.832]). However, both models have large NDE (>0.05>0.05) and NIE (>0.14>0.14). Causally fair resampling maintains near-equal performance to the baseline model (0.809[0.794, 0.825]0.809\ [0.794,\ 0.825]) while eliminating the NDE; however, the NIE is still relatively large (0.111[0.095, 0.115]-0.111\ [0.095,\ 0.115]). Greedy feature selection also eliminates the NDE and improves the NIE beyond the baseline and causally fair resampling (0.074[0.069, 0.087]0.074\ [0.069,\ 0.087]); however, there is a $6.8\%$ drop in AUROC (0.757[0.738, 0.777]0.757\ [0.738,\ 0.777]). If we allow model performance to drop even further, the ”unbiased” feature selection intervention (AUROC: 0.695[0.674, 0.716]0.695\ [0.674,\ 0.716]) reduces the NIE even further (0.027[0.021, 0.046]0.027\ [0.021,\ 0.046]) while eliminating the NDE. Through this, we observe that it is possible to eliminate the NDE without sacrificing model performance, but there is a clear tradeoff between reducing the NIE and maintaining model performance (Appendix D).
5.5.2. Systemic Lupus Erythematosus (SLE)
@@
As shown in Figure 4B, the demographic-unaware model is associated with the best AUROC (0.726[0.662, 0.794]0.726\ [0.662,\ 0.794]). However, the NDE (0.089[0.094,0.077]-0.089\ [-0.094,\ -0.077]) is significantly higher compared to the data and the baseline model. The next-best model performances (outperforming even the baseline model) are associated with causally fair resampling (0.699[0.638, 0.759]0.699\ [0.638,\ 0.759]) and selection of ”unbiased” features (0.680[0.613,0.741]0.680\ [0.613,0.741]). The causally fair resampling reduces, but does not eliminate the NDE (0.030[0.044,0.030]-0.030\ [-0.044,\ -0.030]) and ”unbiased” feature selection eliminates the NDE (Appendix E).
5.5.3. Type 2 Diabetes Mellitus (T2DM)
@@
The highest performing model that eliminates both the NDE (Figure 4C) and NIE (Figure 4F) is the one trained using path-specific in-processing. The AUROC of this model (0.689[0.671,0.709]0.689\ [0.671,0.709]) represents a $7.3\%$ drop in performance relative to the highest-performing model, which uses the causally fair resampling intervention (AUROC: 0.743[0.728,0.760]0.743\ [0.728,0.760]). However, causally fair resampling maintains a significant NDE (0.020[0.006, 0.034]0.020\ [0.006,\ 0.034]) or the NIE (0.041[0.034, 0.078]0.041\ [0.034,\ 0.078]). Thus, we observe a tradeoff between model performance and model bias along both the NDE and NIE (Appendix F).
5.5.4. Schizophrenia (SCZ)
@@
As shown in Figure 4D and 4G, the baseline model is associated with the best performance (AUROC: 0.759[0.748, 0.770]0.759\ [0.748,\ 0.770]) but larger NDE and SE compared to the data (NDE: 0.120[0.104, 0.134]0.120\ [0.104,\ 0.134]; SE: 0.069[0.073,0.061]-0.069\ [-0.073,\ -0.061]). Greedy feature selection maintains the performance of the baseline model (AUROC: 0.758[0.747, 0.768]0.758\ [0.747,\ 0.768]) while eliminating the NDE and maintaining a similar SE to the baseline model (0.063[0.065, 0.049]-0.063\ [-0.065,\ 0.049]). If we allow for deterioration in performance, we find that ”unbiased” feature selection (AUROC: 0.698[0.687, 0.710]0.698\ [0.687,\ 0.710]) eliminates the NDE and reduces the SE to 0.042[0.044,0.037]-0.042\ [-0.044,-0.037]. Thus, while we do not observe a tradeoff between performance and NDE, we do observe a tradeoff between performance and SE (Appendix G).
6. Discussion
@@
Causal effect estimation in high dimensions requires dimensionality reduction via outcome-driven feature selection. The underpinning of this fairness pipeline is the ability to robustly estimate path-specific causal effects in high dimensions of $W$, as the dimensionality of the clinical representation ($W$) may vary from hundreds (AMI, SLE, T2DM) to over $1000$ (SCZ). We leverage state-of-the-art doubly robust estimators, which allow for incorrect specification of either the propensity models or the outcome models, and find that for high-dimensional $W$, estimation error is significantly reduced by selecting features based on their predictive power with respect to $Y$, consistent with mediation analyses in high-throughput data (Jérolon et al., 2024; Chén et al., 2018). In the case where $Z$ were to be a high-dimensional vector, approaches related to high-dimensional propensity analysis could help ensure robust estimates (Brookhart et al., 2006; Schneeweiss et al., 2009; Zhang et al., 2022).
@@
Consider causal and feature selection-based fairness interventions for reducing path-specific causal effects. We test interventions from different types of fairness literature, including group fairness (equalized odds (Hardt et al., 2016)), individual fairness (learning fair representations (Zemel et al., 2013)), and causal fairness (path-specific inprocessing (Plečko and Bareinboim, 2024), causal resampling (Nabi et al., 2019)). Additionally, we test several naive feature selection approaches not designed to target any specific type of fairness (fairness through unawareness (Höltgen and Oliver, 2025), selection of unbiased features, greedy feature selection). No universally optimal approach exists. Interestingly, we find that causal methods do not always have the best efficacy; relatively naive feature selection approaches, including greedy feature selection and selecting ”unbiased” features, can significantly improve causal fairness. We hypothesize that naive feature selection may be a feasible approach in observational health data due to some interdependence between features: even if the ”most biased” features are removed from the model, the patterns that underlie a person’s health are still learnable from other features in the data. We additionally find that non-causal algorithmic fairness interventions (e.g.. methods that target individual (Zemel et al., 2013) and group (Hardt et al., 2016) fairness objectives) are unable to maintain reasonable model performance when targeting the systemic disparities prevalent in healthcare. These results illustrate the importance of implementing the full pipeline (Figure 2) and trying both causal and non-causal interventions to mitigate path-specific causal effects.
@@ 7. Conclusion
In this work, we establish a pipeline for operationalizing path-specific causal fairness in clinical prediction tasks. We define a generalizable mapping of tabular observational health data onto the structural fairness model, and demonstrate how to construct this mapping in the context of known health disparities. We illustrate the importance of defining model fairness through causal pathways that replicate real-world disparity and quantify the extent to which learned models exacerbate these existing biases. By characterizing a more comprehensive fairness-accuracy tradeoff that examines multiple fairness pathways alongside accuracy, we illustrate the importance of context-specific prioritization and enforcement of causal pathways over which a model should be ”fair”. While our work focuses on the algorithms in the healthcare setting, several components of the pipeline (robust estimation of causal effects, selection of causal pathways to target for bias mitigation, trial of causal and non-causal fairness interventions) are applicable to fair ML outside of healthcare.
8. Endmatter
@@
Ethical Considerations: This study was approved by the relevant institutional review board (IRB). Informed consent was waived by the IRB due to low risk to subject welfare and logistical infeasibility of contacting millions of patients from a de-identified database.

### v2

@@ 5.4.2 Systemic Lupus Erythematosus (SLE)
When examining the impact of gender on SLE, the NIE and SE are insignificant, and the NDE is small (0.002[0.002,0.001]-0.002\ [-0.002,-0.001]). The NDE and NIE increase significantly in the baseline model (NDE: 0.038[0.047,0.025]-0.038\ [-0.047,\ -0.025]; NIE: 0.012[0.021,0.008]-0.012\ [-0.021,\ -0.008]), with the NDE increasing more than the NIE. The negative NDE indicates that women are more likely to be diagnosed than men, consistent with the low diagnosis rates of SLE in men (Tan et al., 2012; Simard et al., 2022). The NIE suggests that individuals with a "male" distribution are more likely to be diagnosed with SLE than individuals with data drawn from a "female" distribution. Due to direct bias against men when diagnosing SLE and the larger size of the NDE in both the data and the baseline model, we focus on reducing the NDE for SLE.
5.4.3 Type 2 Diabetes Mellitus (T2DM)
@@
In the T2DM cohort, the dataset reflects a significant NDE (0.014[0.002,0.016]0.014\ [0.002,0.016]) that indicates overdiagnosis of T2DM in Black individuals, consistent with epidemiological literature (Deng et al., 2025; Hackl et al., 2024). The NDE is exacerbated in the baseline model (0.033[0.027,0.134]0.033\ [0.027,0.134]), which also worsens the NIE (0.059[0.053,0.107]0.059\ [0.053,0.107]). The NIE suggests that an individual with clinical features drawn from a "white" population distribution is more likely to be diagnosed with T2DM; this is consistent with known algorithmic biases (Cronjé et al., 2023). The SE remains insignificant in the data and baseline model. Thus, we focus on jointly reducing the NDE and NIE for T2DM risk prediction.
5.4.4 Schizophrenia (SCZ)
@@
The NDE, NIE, and SE are all similar in magnitude (NDE: 0.018[0.005, 0.031]0.018\ [0.005,\ 0.031]; NIE: 0.017[0.003,0.027]0.017\ [0.003,0.027]; SE: 0.016[0.017,0.012]-0.016\ [-0.017,\ -0.012]) in the data. The positive NDE is reflective of overdiagnosis of SCZ in Black patients compared to white patients, while the positive NIE is reflective of the fact that an individual with clinical features drawn from a "white" population distribution is more likely to be diagnosed with SCZ. The negative SE suggests that the confounder variables in $Z$ (gender, healthcare utilization) may play a larger role in diagnosis for Black patients than white patients. Of these three effects, the NDE reflects known overdiagnosis of SCZ in Black individuals (Trierweiler et al., 2000), while the NIE and SE may reflect broader disparities in quality of care (Fiscella and Sanders, 2016) and healthcare utilization Manuel (2018), respectively. In the baseline model, we observe a significant exacerbation in NDE (Model: 0.120[0.104, 0.134]0.120\ [0.104,\ 0.134]) and SE (Model: 0.069[0.073,0.061]-0.069\ [-0.073,-0.061]). The NIE remains similar the data (0.015[0, 0.039]0.015\ [0,\ 0.039]). These causal effects suggest that the model is more reliant on race than a clinician might be (NDE), and that the disparate weight on confounding variables (healthcare utilization, gender) is more pronounced in the model than in clinical practice (SE). Based on the tenfold increase in NDE between the raw data and the baseline model and the known problems with overdiagnosis of schizophrenia in Black individuals, we prioritize reducing the NDE in the SCZ risk prediction case. Additionally, we treat reduction of SE as a secondary target, as the effect may be due to disparities in healthcare utilization.
5.5 Perform and evaluate fairness interventions
@@
We discuss the efficacy and impact of all interventions outlined in Section 4.5 for each clinical prediction task. We present the AUROC-effect tradeoff for each of the "target" effects (Figure 8).
@@
Figure 7: Tradeoffs between model performance and target path-specific effects across all tasks. We plot the target path-specific effects for each intervention (NDE, NIE, and/or SE) against the model performance (AUROC) for each of the "target" effects (Section 4.4). (A) and (E) show the NDE and NIE for AMI; (B) shows the NDE for SLE; (C) and (F) show the NDE and NIE for T2DM; (D) and (G) show the NDE and SE for SCZ.
@@
Figure 8: 4 columns by 2 rows of scatter plots. Points are colored differently based on the fairness intervention (included in the legend) and have associated error bars. The x-axis is the path-specific causal effect and the y-axis shows AUROC.
5.5.1 Acute Myocardial Infarction (AMI)
@@
As shown in Figure 8A and 8E, the baseline model and demographic-unaware model are associated with the highest model performance (Baseline AUROC: 0.812[0.796, 0.825]0.812\ [0.796,\ 0.825]; demographic unawareness: 0.814[0.799, 0.832]0.814\ [0.799,\ 0.832]). However, both models have large NDE (>0.05>0.05) and NIE (>0.14>0.14). Causally fair resampling maintains near-equal performance to the baseline model (0.809[0.794, 0.825]0.809\ [0.794,\ 0.825]) while eliminating the NDE; however, the NIE is still relatively large (0.111[0.095, 0.115]-0.111\ [0.095,\ 0.115]). Greedy feature selection also eliminates the NDE and improves the NIE beyond the baseline and causally fair resampling (0.074[0.069, 0.087]0.074\ [0.069,\ 0.087]); however, there is a $6.8\%$ drop in AUROC (0.757[0.738, 0.777]0.757\ [0.738,\ 0.777]). If we allow model performance to drop even further, the "unbiased" feature selection intervention (AUROC: 0.695[0.674, 0.716]0.695\ [0.674,\ 0.716]) reduces the NIE even further (0.027[0.021, 0.046]0.027\ [0.021,\ 0.046]) while eliminating the NDE. Through this, we observe that it is possible to eliminate the NDE without sacrificing model performance, but there is a clear tradeoff between reducing the NIE and maintaining model performance (Appendix 12).
5.5.2 Systemic Lupus Erythematosus (SLE)
@@
As shown in Figure 8B, the demographic-unaware model is associated with the best AUROC (0.726[0.662, 0.794]0.726\ [0.662,\ 0.794]). However, the NDE (0.089[0.094,0.077]-0.089\ [-0.094,\ -0.077]) is significantly higher compared to the data and the baseline model. The next-best model performances (outperforming even the baseline model) are associated with causally fair resampling (0.699[0.638, 0.759]0.699\ [0.638,\ 0.759]) and selection of "unbiased" features (0.680[0.613,0.741]0.680\ [0.613,0.741]). The causally fair resampling reduces, but does not eliminate the NDE (0.030[0.044,0.030]-0.030\ [-0.044,\ -0.030]) and "unbiased" feature selection eliminates the NDE (Appendix 13).
5.5.3 Type 2 Diabetes Mellitus (T2DM)
@@
The highest performing model that eliminates both the NDE (Figure 8C) and NIE (Figure 8F) is the one trained using path-specific in-processing. The AUROC of this model (0.689[0.671,0.709]0.689\ [0.671,0.709]) represents a $7.3\%$ drop in performance relative to the highest-performing model, which uses the causally fair resampling intervention (AUROC: 0.743[0.728,0.760]0.743\ [0.728,0.760]). However, causally fair resampling maintains a significant NDE (0.020[0.006, 0.034]0.020\ [0.006,\ 0.034]) or the NIE (0.041[0.034, 0.078]0.041\ [0.034,\ 0.078]). Thus, we observe a tradeoff between model performance and model bias along both the NDE and NIE (Appendix 14).
5.5.4 Schizophrenia (SCZ)
@@
As shown in Figure 8D and 8G, the baseline model is associated with the best performance (AUROC: 0.759[0.748, 0.770]0.759\ [0.748,\ 0.770]) but larger NDE and SE compared to the data (NDE: 0.120[0.104, 0.134]0.120\ [0.104,\ 0.134]; SE: 0.069[0.073,0.061]-0.069\ [-0.073,\ -0.061]). Greedy feature selection maintains the performance of the baseline model (AUROC: 0.758[0.747, 0.768]0.758\ [0.747,\ 0.768]) while eliminating the NDE and maintaining a similar SE to the baseline model (0.063[0.065, 0.049]-0.063\ [-0.065,\ 0.049]). If we allow for deterioration in performance, we find that "unbiased" feature selection (AUROC: 0.698[0.687, 0.710]0.698\ [0.687,\ 0.710]) eliminates the NDE and reduces the SE to 0.042[0.044,0.037]-0.042\ [-0.044,-0.037]. Thus, while we do not observe a tradeoff between performance and NDE, we do observe a tradeoff between performance and SE (Appendix 15).
6 Discussion
@@
Causal effect estimation in high dimensions requires dimensionality reduction via outcome-driven feature selection. The underpinning of this fairness pipeline is the ability to robustly estimate path-specific causal effects in high dimensions of $W$, as the dimensionality of the clinical representation ($W$) may vary from hundreds (AMI, SLE, T2DM) to over $1000$ (SCZ). We leverage state-of-the-art doubly robust estimators, which allow for incorrect specification of either the propensity models or the outcome models, and find that for high-dimensional $W$, estimation error is significantly reduced by selecting features based on their predictive power with respect to $Y$, consistent with mediation analyses in high-throughput data (Jérolon et al., 2024; Chén et al., 2018). In the case where $Z$ were to be a high-dimensional vector, approaches related to high-dimensional propensity analysis could help ensure robust estimates Brookhart et al. (2006); Schneeweiss et al. (2009); Zhang et al. (2022).
@@
Consider causal and feature selection-based fairness interventions for reducing path-specific causal effects. We test interventions from different types of fairness literature, including group fairness (equalized odds (Hardt et al., 2016)), individual fairness (learning fair representations (Zemel et al., 2013)), and causal fairness (path-specific inprocessing (Plečko and Bareinboim, 2024), causal resampling (Nabi et al., 2019)). Additionally, we test several naive feature selection approaches not designed to target any specific type of fairness (fairness through unawareness (Höltgen and Oliver, 2025), selection of unbiased features, greedy feature selection). No universally optimal approach exists. Interestingly, we find that causal methods do not always have the best efficacy; relatively naive feature selection approaches, including greedy feature selection and selecting "unbiased" features, can significantly improve causal fairness. We hypothesize that naive feature selection may be a feasible approach in observational health data due to some interdependence between features: even if the "most biased" features are removed from the model, the patterns that underlie a person’s health are still learnable from other features in the data. We additionally find that non-causal algorithmic fairness interventions (e.g.. methods that target individual (Zemel et al., 2013) and group (Hardt et al., 2016) fairness objectives) are unable to maintain reasonable model performance when targeting the systemic disparities prevalent in healthcare. These results illustrate the importance of implementing the full pipeline (Figure 4) and trying both causal and non-causal interventions to mitigate path-specific causal effects.
@@ 7 Conclusion
In this work, we establish a pipeline for operationalizing path-specific causal fairness in clinical prediction tasks. We define a generalizable mapping of tabular observational health data onto the structural fairness model, and demonstrate how to construct this mapping in the context of known health disparities. We illustrate the importance of defining model fairness through causal pathways that replicate real-world disparity and quantify the extent to which learned models exacerbate these existing biases. By characterizing a more comprehensive fairness-accuracy tradeoff that examines multiple fairness pathways alongside accuracy, we illustrate the importance of context-specific prioritization and enforcement of causal pathways over which a model should be "fair". While our work focuses on the algorithms in the healthcare setting, several components of the pipeline (robust estimation of causal effects, selection of causal pathways to target for bias mitigation, trial of causal and non-causal fairness interventions) are applicable to fair ML outside of healthcare.
8 Endmatter
Ethical Considerations:
@@ 8 Endmatter
This study was approved by the relevant institutional review board (IRB). Informed consent was waived by the IRB due to low risk to subject welfare and logistical infeasibility of contacting millions of patients from a de-identified database.
Generative AI usage statement: